# Subjective-Aligned Dataset and Metric for Text-to-Video Quality Assessment

### v3

@@ Abstract.
With the rapid development of generative models, AI-Generated Content (AIGC) has exponentially increased in daily lives. Among them, Text-to-Video (T2V) generation has received widespread attention. Though many T2V models have been released for generating high perceptual quality videos, there is still lack of a method to evaluate the quality of these videos quantitatively. To solve this issue, we establish the largest-scale Text-to-Video Quality Assessment DataBase (T2VQA-DB) to date. The dataset is composed of 10,000 videos generated by 9 different T2V models, along with each video’s corresponding mean opinion score. Based on T2VQA-DB, we propose a novel transformer-based model for subjective-aligned Text-to-Video Quality Assessment (T2VQA). The model extracts features from text-video alignment and video fidelity perspectives, then it leverages the ability of a large language model to give the prediction score. Experimental results show that T2VQA outperforms existing T2V metrics and SOTA video quality assessment models. Quantitative analysis indicates that T2VQA is capable of giving subjective-align predictions, validating its effectiveness. The dataset and code will be released at https://github.com/QMME/T2VQA.
@@ 1. Introduction
Video generation, or video synthesis, has been fully developed in the past few years. Text-to-Video (T2V) generation is one of the most studied fields, where a user provides a text description as the guidance for video generation. With the thriving of diffusion-based models, high-fidelity videos can be generated. However, the quality of text-generated videos is diverse which affects the experience quality of subjects. Therefore, a subjective-aligned quality assessment method for them is needed. Unfortunately, existing Video Quality Assessment (VQA) models are unable to accomplish the task well. On the one hand, distortions brought by the T2V generation models, such as the jitter effect, irrational objects, etc, are different from distortions in natural videos. On the other hand, traditional VQA models do not take text-video alignment into consideration, which is a significant evaluation perspective for text-generated videos.
@@
To facilitate the development of a more comprehensive and accurate metric, we establish the largest-scale subjective T2V dataset to date, named Text-to-Video Quality Assessment DataBase (T2VQA-DB). The dataset contains 10,000 videos generated by 9 different T2V models using 1,000 text prompts. We also collect each video’s Mean Opinion Score (MOS) by conducting a subjective experiment, where 27 subjects score the overall quality of the generated videos. Fig. 2 shows video examples from T2VQA-DB. We anticipate that the T2VQA-DB will benefit the training and testing of subsequent models.
Based on T2VQA-DB, we propose a novel model equipped with multi-modality foundation models for better Text-to-Video Quality Assessment (T2VQA). The model utilizes BLIP (Li et al., 2022a) and Video Swin Transformer (Swin-T) (Liu et al., 2022) to extract features from text-video alignment and video fidelity perspectives respectively. The features are fused through a cross-attention module, then they are fed into a frozen Large Language Model (LLM) to regress the predicted score. We train and test T2VQA as well as other VQA models on T2VQA-DB. Experimental results show that T2VQA outperforms existing T2V generation metrics and state-of-the-art VQA models, validating its effectiveness in measuring the perceptual quality of text-generated videos.
@@
The proposed T2VQA outperforms existing T2V generation metrics and SOTA VQA models on T2VQA-DB and Sora (Brooks et al., 2024) videos, indicating the effectiveness of T2VQA. Qualitative experiments show that T2VQA can benefit in measuring the performance of T2V generation algorithms, giving it practical application prospects.
@@ 2.1. Text-to-video Dataset
To the best of our knowledge, only a few T2V datasets have been proposed. They mainly have the following two issues: (1) Insufficient scale: Chivileva et al. (Chivileva et al., 2023) proposes a dataset with 1,005 videos generated by 5 T2V models following (Li et al., 2023b). EvalCrafter (Liu et al., 2023) builds a dataset using 500 prompts and 7 T2V models, resulting in 3,500 videos in total. Similarly, FETV (Liu et al., 2024) is composed of 2,476 videos generated by 619 prompts, and 4 T2V models. VBench (Huang et al., 2024) has a larger scale with in total of 1.7k prompts and 4 T2V models. Such scales are not sufficient for the training of deep learning-based models, and cannot comprehensively represent current T2V algorithms. (2) Limited human annotation: ITU-standard (ITU-R, 2002) requires at least 15 human annotators for subjective study. The dataset proposed by Chivileva et al. (Chivileva et al., 2023) is the only one that meets the standard with 24 annotators involved. Both EvalCrafter and FETV only have 3 users for annotation. VBench (Huang et al., 2024) does not specify the number of human annotators explicitly.
Name | Videos | Prompts | Models | Annotators |
Chivileva’s (Chivileva et al., 2023) | 1,005 | 201 | 5 | 24 |
EvalCrafter (Liu et al., 2023) | 3,500 | 500 | 7 | 3 |
VBench (Huang et al., 2024) | 6,984 | 1,746 | 4 | - |
FETV (Liu et al., 2024) | 2,476 | 619 | 4 | 3 |
T2VQA-DB (ours) | 10,000 | 1,000 | 9 | 27 |
@@ 2.2. Metrics for T2V Generation
Though many works targeting the evaluation of text-generated images have been proposed (Xu et al., 2024; Kirstain et al., 2024; Wu et al., 2023c; Li et al., 2023b, 2024; Zhang et al., 2023a), only a few metrics tailored for text-generated video evaluation have been proposed. Among them, ViCLIP (Wang et al., 2023b) is a CLIP-based metric for measuring text-video alignment. Chivileva et al. (Chivileva et al., 2023) proposes an ensemble video quality metric that integrates text similarity and naturalness. EvalCrafter (Liu et al., 2023) and VBench (Huang et al., 2024) build benchmarks to evaluate text-generated video from 18 and 16 objective metrics respectively. FETV (Liu et al., 2024) and T2V-Score (Wu et al., 2024a) both propose separate metrics for text-video alignment and video quality, without an overall perceptual score for text-generated videos. There is still lack of a simple and effective metric to evaluate the quality of text-generated videos accurately.
Besides the aforementioned metrics, VQA models can also be used for the evaluation of text-generated videos. BVQA (Li et al., 2022b) transfers knowledge from Image Quality Assessment (IQA) databases and then trains on the target VQA database. SimpleVQA (Sun et al., 2022) extracts spatial and motion features to regress to the final score. FAST-VQA (Wu et al., 2022) proposes “fragments” as a novel sampling strategy and the Fragment Attention Network (FANet) to accommodate fragments as inputs. (Gao et al., 2023; Dong et al., 2023; Kou et al., 2023; Zhang et al., 2023b) are works for the evaluation of enhanced videos and digital humans. DOVER (Wu et al., 2023e) proposes to view the quality assessment problem from the technical perspective and the aesthetic perspective, while BVQI (Wu et al., 2023b) and MaxVQA (Wu et al., 2023d) integrates text prompts (e.g., good, bad) into VQA. With the development of LLMs and Multi-modal Large Language Models (MLLM), researchers have started to leverage the advantages of MLLMs to solve VQA problems, as they have trained on massive data. Q-Bench (Wu et al., 2023g) proves that MLLMs can address preliminary low-level visual tasks. Q-Align (Wu et al., 2023f) proposes to train MLLMs using text-defined levels (e.g., fine, poor, excellent) and achieves SOTA results on both IQA and VQA tasks. Though many VQA models have been proposed, they are originally designed for natural videos and do not consider text-video alignment.
@@ 3. Subjective-Aligned Text-to-Video Dataset
As shown in Tab. 1, the existing T2V datasets have a relatively small number of videos. The small scale is not sufficient to represent the diverse performance of T2V generation models, resulting in unreliable quality assessment metrics. Consequently, we propose a Text-to-Video Quality Assessment DataBase, named T2VQA-DB, including 10,000 videos generated by 9 different T2V models. 1,000 prompts are used and 27 subjects are invited to obtain the MOS of each video. In this section, we will describe the establishment of T2VQA-DB and the subjective experiment conducted on it.
3.1. Prompt Selection
@@
3.2. Video Generation
We use in total 9 different models for video generation, including Text2Video-Zero (Khachatryan et al., 2023), AnimateDiff (Guo et al., 2023), Tune-a-video (Wu et al., 2023a), VidRD (Gu et al., 2023), VideoFusion (Luo et al., 2023), ModelScope (Wang et al., 2023c), LVDM (He et al., 2022), Show-1 (Zhang et al., 2023d), and LaVie (Wang et al., 2023a). For Tune-a-video, we utilize two different pre-trained weights, resulting in a total of 10 models for generation. Compared to other T2V datasets, we utilize current advanced T2V generation models as much as possible, making T2VQA-DB more representative. Since the default resolution, video length, and frame rate are different in each model, we unify the video format as , 16 frames, and 4fps. We end up generating 10,000 text-generated videos.
@@ 3.3. Subjective Study
where and refer to the index of videos and subjects. is the raw score, and , . is the number of videos scored by each subject. is the number of scores on one video. is the rescaling function, converting the distribution of Z-scores into a mean of 50 and a standard deviation of 16.6. Fig. 3b shows the distributions of raw MOS and Z-score MOS.
@@ 3.4. Dataset Analysis
We also conduct comprehensive experiments and analysis on T2VQA-DB. We first investigate each model’s performance in T2VQA-DB. The visualization of the Z-score MOS distributions of 10 models is shown in Fig. 3c. LaVie has the highest average MOS of 66.9, while two Tune-a-Video models have the lowest of 39.1 and 39.9. The reason for the poor performance of Tune-a-Video is mainly the low inter-frame consistency, as shown in Fig. 7c.
Based on the prompt contents, we classify the collected prompts into 6 categories, including nature, human, artificial, animal, object, and abstract. The ones that cannot be categorized into the 6 classes are labeled as “others”. After classification, we have 327 prompts for nature, 119 for human, 196 for artificial, 121 for animal, 66 for object, 135 for abstract, and 36 for others. Subsequently, we compare the models’ performance over different prompt types. As shown in Fig. 4, LaVie outperforms the other models on all types of prompts. Tune-a-video has the worst performance, which is consistent with the analysis in Fig. 3c. LVDM has the second-worst performance except in the “others” type. The performance of the rest models has negligible differences. Fig. 4 also shows prompts classified as human have the worst performance in all models. The reason could be that human faces and actions require more sophisticated modeling compared with other categories. Some models show a preference for object and nature types of prompts to a small extent. The models do not show obvious preferences for other types of prompts.
4. Subjective-Aligned Text-to-video Metric
@@ 4.1. Text-video Alignment Encoder
Text-video alignment refers to the conformity between the video content and the text description. CLIP (Radford et al., 2021) and BLIP (Li et al., 2022a) are models that have a strong ability for zero-shot text-image matching. (Wu et al., 2021) first proposes CLIPSim, which uses CLIP to calculate the similarities between text and each frame of the video and then take the average value. However, former works (Chivileva et al., 2023; Huang et al., 2024; Liu et al., 2024, 2023) and results in Tab. 2 show that simply using CLIP or BLIP has a low correlation with the authentic subjective scores. Following works of MLLMs such as BLIP-2 (Li et al., 2023a), InstructBLIP (Dai et al., 2023), and mPLUG-Owl2 (Ye et al., 2023), we use the pre-trained BLIP image encoder as the video frame encoder. We freeze the weights of the image encoder that encodes each frame separately. We use the BLIP text encoder as the alignment encoder. The alignment encoder takes the encoded text and video frame as inputs. They interact through the cross-attention module and the encoder eventually outputs a feature representing the text-frame similarity. We concatenate features from each text-frame pair. Given a set of video frames and the text prompt , we have:

### v5

@@ Abstract.
With the rapid development of generative models, AI-Generated Content (AIGC) has exponentially increased in daily lives. Among them, Text-to-Video (T2V) generation has received widespread attention. Though many T2V models have been released for generating high perceptual quality videos, there is still lack of a method to evaluate the quality of these videos quantitatively. To solve this issue, we establish the largest-scale Text-to-Video Quality Assessment DataBase (T2VQA-DB) to date. The dataset is composed of 10,000 videos generated by 9 different T2V models, along with each video’s corresponding mean opinion score. Based on T2VQA-DB, we propose a novel transformer-based model for subjective-aligned Text-to-Video Quality Assessment (T2VQA). The model extracts features from text-video alignment and video fidelity perspectives, then it leverages the ability of a large language model to give the prediction score. Experimental results show that T2VQA outperforms existing T2V metrics and SOTA video quality assessment models. Quantitative analysis indicates that T2VQA is capable of giving subjective-align predictions, validating its effectiveness. The dataset and code are available at https://github.com/QMME/T2VQA.
@@ 1. Introduction
Video generation, or video synthesis, has been fully developed in the past few years. Text-to-Video (T2V) generation is one of the most studied fields, where a user provides a text description as the guidance for video generation. With the thriving of diffusion-based models, high-fidelity videos can be generated. However, the quality of text-generated videos is diverse which affects the experience quality of subjects. Therefore, a subjective-aligned quality assessment method for them is needed. Unfortunately, existing Video Quality Assessment (VQA) models are unable to accomplish the task well. On the one hand, distortions brought by T2V generation models, such as jitter effect, irrational objects, are different from distortions in natural videos. On the other hand, traditional VQA models do not take text-video alignment into consideration, which is a significant evaluation perspective for text-generated videos.
@@
To facilitate the development of a more comprehensive and accurate metric, we establish the largest-scale subjective T2V dataset to date, named Text-to-Video Quality Assessment DataBase (T2VQA-DB). The dataset contains 10,000 videos generated by 9 representative T2V models using 1,000 text prompts. We also collect each video’s Mean Opinion Score (MOS) by conducting a subjective experiment, where 27 subjects score the overall quality of the generated videos. Fig. 2 shows video frames generated by the 9 T2V models from the prompt “Sunset on the sea”. And we give more detailed overviews of three examples. We anticipate that the T2VQA-DB will benefit the development of subsequent models.
Based on T2VQA-DB, we propose a novel model equipped with multi-modality foundation models for better Text-to-Video Quality Assessment (T2VQA). The model utilizes BLIP (Li et al., 2022a) and Video Swin Transformer (Swin-T) (Liu et al., 2022) to extract features from text-video alignment and video fidelity perspectives respectively. The features are fused through a cross-attention module, then they are fed into a frozen Large Language Model (LLM) to regress the predicted score. We train and test T2VQA as well as other VQA models on T2VQA-DB. Experimental results show that T2VQA outperforms existing T2V generation metrics and state-of-the-art VQA models, validating its effectiveness in measuring the perceptual quality of text-generated videos. Fig. 1 shows the overview of the proposed T2VQA-DB and T2VQA.
@@
The proposed T2VQA outperforms existing T2V generation metrics and SOTA VQA models on T2VQA-DB and other T2V datasets, indicating the effectiveness of T2VQA. Qualitative experiments show that T2VQA can benefit in measuring the performance of T2V generation algorithms, giving it practical application prospects.
| Name | Videos | Prompts | Models | Annotators |
| Chivileva’s (Chivileva et al., 2023) | 1,005 | 201 | 5 | 24 |
| ECTV (Liu et al., 2023) | 5,600 | 700 | 8 | 7 |
| VBench (Huang et al., 2024) | 6,984 | 1,746 | 4 | - |
| FETV (Liu et al., 2024) | 2,476 | 619 | 4 | 3 |
| TVGE (Wu et al., 2024a) | 2,543 | 512 | 5 | 10 |
| T2VQA-DB (ours) | 10,000 | 1,000 | 9 | 27 |
@@ 2.1. Text-to-video Dataset
To the best of our knowledge, only a few T2V datasets have been proposed. They mainly have the following two issues: (1) Insufficient scale: Chivileva et al. (Chivileva et al., 2023) propose a dataset with 1,005 videos generated by 5 T2V models following (Li et al., 2023b). EvalCrafter (Liu et al., 2023) builds ECTV using 700 prompts and 8 T2V models, resulting in 5,600 videos in total. Similarly, FETV (Liu et al., 2024) is composed of 2,476 videos generated by 619 prompts, and 4 T2V models. VBench (Huang et al., 2024) has a larger scale with in total of 1.7k prompts and 4 T2V models. TVGE (Wu et al., 2024a) uses the prompts from ECTV and uses 5 T2V model for video generation. Such scales are not sufficient for the training of deep learning-based models, and cannot represent current T2V algorithms. (2) Limited human annotation: ITU-standard (ITU-R, 2002) requires at least 15 human annotators for subjective study. The dataset proposed by Chivileva et al.is the only one that meets the standard. ECTV and FETV only have 3 users for annotation. VBench (Huang et al., 2024) does not specify the number of human annotators.
@@ 2.2. Metrics for T2V Generation
Though many works targeting the evaluation of text-generated images have been proposed (Xu et al., 2024; Kirstain et al., 2024; Wu et al., 2023c; Li et al., 2023b, 2024b; Zhang et al., 2023a; Li et al., 2024a; Yang et al., 2024), there are only a few metrics tailored for text-generated video evaluation have been proposed. Among them, ViCLIP (Wang et al., 2023b) is a CLIP-based metric for measuring text-video alignment. Chivileva et al. (Chivileva et al., 2023) proposes an ensemble video quality metric that integrates text similarity and naturalness. EvalCrafter (Liu et al., 2023) and VBench (Huang et al., 2024) build benchmarks to evaluate text-generated video from 18 and 16 objective metrics respectively. FETV (Liu et al., 2024) and T2V-Score (Wu et al., 2024a) both propose separate metrics for text-video alignment and video quality, without an overall perceptual score for text-generated videos. There is still lack of a simple and effective metric to evaluate the quality of text-generated videos.
Besides the aforementioned metrics, VQA models can also be used for the evaluation of text-generated videos. BVQA (Li et al., 2022b) transfers knowledge from Image Quality Assessment (IQA) databases and then trains on the target VQA database, e.g., KoNVid-1k(Hosu et al., 2017). SimpleVQA (Sun et al., 2022) extracts spatial and motion features to regress to the final score. FAST-VQA (Wu et al., 2022) proposes “fragments” as a novel sampling strategy and the fragment attention network to accommodate fragments as inputs. (Gao et al., 2023; Dong et al., 2023; Kou et al., 2023; Zhang et al., 2023b, 2024b; Zhou et al., 2024a, b) are works for the evaluation of enhanced videos and digital humans. DOVER (Wu et al., 2023e) proposes to view the quality assessment problem from technical and aesthetic perspectives, while BVQI (Wu et al., 2023b) and MaxVQA (Wu et al., 2023d) integrate text prompts (e.g., good, bad) into VQA. With the development of Multi-modal Large Language Models (MLLMs), researchers have started to leverage MLLMs to solve VQA problems, as they have been trained on massive data. Q-Bench (Wu et al., 2023g) proves that MLLMs can address preliminary low-level visual tasks. Q-Align (Wu et al., 2023f) proposes to train a MLLM using text-defined levels (e.g., fine, poor) and achieves SOTA results on IQA and VQA tasks. Though many VQA models have been proposed, they are originally designed for natural videos and do not consider text-video alignment.
@@ 3. Subjective-Aligned Text-to-Video Dataset
Tab. 1 shows most existing T2V datasets have relatively small number of videos, which is not sufficient to represent the diverse performance of T2V generation models. Consequently, we propose a Text-to-Video Quality Assessment DataBase, named T2VQA-DB, including 10,000 videos generated by 9 different T2V models. 1,000 prompts are used and 27 subjects are invited to obtain the MOS of each video. In this section, we will describe the establishment of T2VQA-DB and the subjective experiment.
3.1. Data Preparation
3.1.1. Prompt Selection.
@@ 3.1. Data Preparation
3.1.2. Video Generation.
We use in total 9 different models for video generation, including Text2Video-Zero (Khachatryan et al., 2023), AnimateDiff (Guo et al., 2023), Tune-a-video (Wu et al., 2023a), VidRD (Gu et al., 2023), VideoFusion (Luo et al., 2023), ModelScope (Wang et al., 2023c), LVDM (He et al., 2022), Show-1 (Zhang et al., 2023d), and LaVie (Wang et al., 2023a). For Tune-a-video, we utilize two different pre-trained weights, resulting in a total of 10 models for generation. Compared to other T2V datasets, we utilize current advanced T2V generation models as much as possible, making T2VQA-DB more representative. Since the default resolution, video length, and frame rate are different in each model, we unify the video format as , 16 frames, and 4fps.
@@ 3.2. Subjective Study
and refer to the index of videos and subjects. is the raw score, and , . is the number of videos scored by each subject. is the number of scores on one video. is the rescaling function, converting the distribution of Z-scores into a mean of 50 and a standard deviation of 16.6. Fig. 3b shows the distributions of raw MOS and Z-score MOS.
@@ 3.3. Dataset Analysis
We conduct comprehensive experiments and analysis on T2VQA-DB. We first investigate each model’s performance in T2VQA-DB. The visualization of the Z-score MOS distributions of 10 models is shown in Fig. 3c. LaVie has the highest average MOS of 66.9, while two Tune-a-Video models have the lowest of 39.1 and 39.9. The reason for the poor performance of Tune-a-Video is mainly the low inter-frame consistency, as shown in Fig. 7c.
Based on the prompt contents, we classify the collected prompts into 6 categories, including nature, human, artificial, animal, object, and abstract. The ones that cannot be categorized into the 6 classes are labeled as “others”. After classification, we have 327 prompts for nature, 119 for human, 196 for artificial, 121 for animal, 66 for object, 135 for abstract, and 36 for others. Subsequently, we compare the models’ performance over different prompt types. As shown in Fig. 4, LaVie outperforms the other models on all types of prompts. Tune-a-video has the worst performance, which is consistent with the analysis in Fig. 3c. LVDM has the second-worst performance except in the “others” type. The performance of the rest models has negligible differences. Fig. 4 also shows prompts labeled as human have the worst performance in all models. The reason could be that human faces and actions require more sophisticated modeling compared with other categories. Some models also show a preference for object and nature types of prompts to a small extent.
| Type | Models | T2VQA-DB Validation | ECTV Testing | TVGE Testing | Sora Testing | ||||||
| SRCC | PLCC | KRCC | RMSE | SRCC | PLCC | SRCC | PLCC | SRCC | PLCC | ||
| zero-shot | CLIPSim (Radford et al., 2021) | 0.1047 | 0.1277 | 0.0702 | 21.683 | 0.3111 | 0.3175 | 0.2765 | 0.3013 | 0.2116 | 0.1538 |
| BLIP (Li et al., 2022a) | 0.1659 | 0.1860 | 0.1112 | 18.373 | 0.3841 | 0.3877 | 0.3437 | 0.3696 | 0.2126 | 0.1038 | |
| ImageReward (Xu et al., 2024) | 0.1875 | 0.2121 | 0.1266 | 18.243 | 0.5192 | 0.5107 | 0.4740 | 0.4879 | 0.0992 | 0.0415 | |
| ViCLIP (Wang et al., 2023b) | 0.1162 | 0.1449 | 0.0781 | 21.655 | 0.4130 | 0.4105 | 0.3429 | 0.3691 | 0.2567 | 0.1844 | |
| UMTScore (Liu et al., 2024) | 0.0676 | 0.0721 | 0.0453 | 22.559 | 0.2406 | 0.2214 | 0.2155 | 0.2199 | 0.2594 | 0.0840 | |
| finetuned | SimpleVQA (Sun et al., 2022) | 0.6275 | 0.6338 | 0.4466 | 11.163 | 0.1959 | 0.1391 | 0.3297 | 0.2219 | 0.0340 | 0.2344 |
| BVQA (Li et al., 2022b) | 0.7390 | 0.7486 | 0.5487 | 15.645 | 0.2929 | 0.2829 | 0.4283 | 0.4229 | 0.4235 | 0.2489 | |
| FAST-VQA (Wu et al., 2022) | 0.7173 | 0.7295 | 0.5303 | 10.595 | 0.3101 | 0.3246 | 0.3335 | 0.3466 | 0.4301 | 0.2369 | |
| DOVER (Wu et al., 2023e) | 0.7609 | 0.7693 | 0.5704 | 9.8072 | 0.3118 | 0.3289 | 0.3943 | 0.4058 | 0.4421 | 0.2689 | |
| Q-Align (Wu et al., 2023f) | 0.7601 | 0.7768 | 0.5860 | 10.9911 | 0.4356 | 0.4401 | 0.5317 | 0.5209 | 0.3109 | 0.1997 | |
| Ours | T2VQA | 0.7965 | 0.8066 | 0.6058 | 9.0221 | 0.5694 | 0.5823 | 0.6898 | 0.6912 | 0.6485 | 0.3124 |
4. Subjective-Aligned T2V Metric
@@ 4.1. Text-video Alignment Encoder
Text-video alignment refers to the conformity between the video content and the text description. CLIP (Radford et al., 2021) and BLIP (Li et al., 2022a) have strong abilities for zero-shot text-image matching. (Wu et al., 2021) proposes CLIPSim, which uses CLIP to calculate the similarities between text and each frame of the video and then take the average value. However, former works (Chivileva et al., 2023; Huang et al., 2024; Liu et al., 2024, 2023) and results in Tab. 2 show that simply using CLIP or BLIP has a low correlation with the authentic subjective scores. Following works of MLLMs (Li et al., 2023a; Dai et al., 2023; Ye et al., 2023), we use the pre-trained BLIP image encoder as the video frame encoder. We freeze the weights of the image encoder that encodes each frame separately. We use the BLIP text encoder as the alignment encoder. The alignment encoder takes the encoded text and video frame as inputs. They interact through the cross-attention module and the encoder eventually outputs a feature representing the text-frame similarity. We concatenate features from each text-frame pair. Given a set of video frames and the text prompt , we have:
where is short for concatenation.